2.5g cane sugar is dissolve in 147.5g of water find the percent by weight of the solution

Grade:12

1 Answers

Arun
5 years ago
Dear student....2.5g + 147.5g = 150g(solution) So atq formula weight of solute/weight of solution×100=2.5×100/150=16.6%